OpenStudy (anonymous):

Identify the type of the equation 4x^2-y^2-16x-4y-4=0 and then change it to standard form.

OpenStudy (aum):

Group the x terms together and complete the square. Group the y terms together and complete the square. Complete these two steps first.

OpenStudy (anonymous):

kk

OpenStudy (aum):

4x^2-y^2-16x-4y-4 = 0 (4x^2 - 16x) - (y^2 + 4y) = 4 4(x^2 - 4x) - (y^2 + 4y) = 4 complete the square of: (x^2 - 4x) complete the square of: (y^2 + 4y)

OpenStudy (anonymous):

got it

OpenStudy (aum):

This equation represents a hyperbola whose standard form looks like this:\[ \frac{(x-h)^2}{a^2} - \frac{(y-k)^2}{b^2} = 1 \]Completing the square will help in putting the equation in standard form.

OpenStudy (anonymous):

How do you complete the square

OpenStudy (aum):

\[ \text{Complete the square:} \\ x^2 - 4x = x^2 - 4x + 4 - 4 = (x-2)^2 - 4 \\ y^2 + 4y = y^2 + 4y + 4 - 4 = (y+2)^2 - 4 \\ \text{Substitute the above two in the earlier equation: } 4(x^2 - 4x) - (y^2 + 4y) = 4 \\ 4{(x-2)^2 - 4} - {(y+2)^2 - 4} = 4 \\ 4(x-2)^2 - 16 - (y+2)^2 + 4 = 4 \\ 4(x-2)^2 - (y+2)^2 = 4 + 16 - 4 \\ 4(x-2)^2 - (y+2)^2 = 16 \\ \text{Divide by 16 to make the right hand side 1:} \\ \frac{(x-2)^2}{4} - \frac{(y+2)^2}{16} = 1 \\ \text{That is the standard form of a hyperbola.} \]